OpenStudy (destinykiara99):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f of x equals four divided by x. and g of x equals four divided by x

OpenStudy (destinykiara99):

@hero @Jaynator495

satellite73 (satellite73):

\[f(g(x))=f(\frac{4}{x})\] is a start

satellite73 (satellite73):

then since \[f(\diamondsuit)=\frac{4}{\diamondsuit}\] you have \[f(\frac{4}{x})=\frac{4}{\frac{4}{x}}\] and then a little algebra

OpenStudy (destinykiara99):

Okay, where do you go from there?? I'm sorry, Pre-Calc is very confusing to me and I don't fully understand the lesson.

OpenStudy (destinykiara99):

@satellite73

satellite73 (satellite73):

multiply by the reciprocal \[\frac{4}{\frac{4}{x}}=4\times \frac{x}{4}=x\]

satellite73 (satellite73):

just like \[\frac{4}{\frac{4}{5}}=4\times \frac{5}{4}=5\]

OpenStudy (destinykiara99):

That makes sense :)

OpenStudy (skullpatrol):

Recall a variable is a symbol used to represent one or more numbers. So that means you can treat them as numbers and use all the operations you have been doing with numbers on them.